# Perspectives on the arithmetic nature of the ratios $\zeta(2n+1)/\pi^{2n+1}$ and $\beta(2n)/\pi^{2n}$

### v2

@@
Abstract: We investigate the values of the Riemann zeta function at odd integers and the Dirichlet beta function at even integers, by collecting several distinct analytic frameworks converging to these values, thus providing a unifying perspective. Beyond analytic interest, these formulas motivate linear independence conjectures which, if established, would imply the irrationality of the quantities $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$.
@@ Introduction
According to Euler’s formula for $\zeta(2n)$ and the expression of $\beta(2n+1)$, many tried to find some rational coefficients $r_{n}$ and $s_{n}$ such that the relations may be satisfied
@@
Given the futility of the extraordinarily numerous attempts to compute these numbers, the solution of the designated problem is generally regarded as impossible; but so far a strong proof of this impossibility is lacking. Only heuristic methods and deep computations by brute force support this impossibility to this day. One may solve the problem by proving that the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ are not multiples of $\pi$. The difficulty of this new reformulation partly stems from the lack of sufficiently many closed-form expressions or identities involving these values that might serve as a starting point for an irrationality or transcendence proof. This paper is devoted to some integral representations of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ through interesting logarithmic, trigonometric and hyperbolic kernels, along with several conjectures that support their non representability as rational multiples of $\pi$.
@@
that has been firstly evaluated to $\dfrac{7\zeta(3)}{8\pi^{2}}$ by Iaroslav V, Blagouchine in 2014 [4]. Therefore the proof of its convergence is omitted here for the sake of conciseness. Notwithstanding, the presence of singularities at its bounds requires to treat it with some care. As consequence, we rewrite the integral as the limit $I=\displaystyle\lim_{\xi\to 0^{+}}\int_{\xi}^{1-\xi}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx$
@@
$I=\displaystyle\lim_{\xi\to 0^{+}}\dfrac{1}{2}\int_{\xi}^{1-\xi}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}dx=\int_{0}^{1}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}\,dx$
The change of variable $x\rightarrow\tan u$ is proceeded in this manner :
@@
$\Rightarrow\displaystyle\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4u}{\ln(\tan u)}\,du=-\dfrac{7\zeta(3)}{\pi^{2}}\quad(1.1)$
@@
$\dfrac{\zeta(5)}{\pi^{4}}=-\dfrac{1}{186}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4x}{\ln(\tan x)}\,dx+\dfrac{1}{124}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 8x}{\ln(\tan x)}\,dx\quad(1.2)$
$\frac{\zeta(7)}{\pi^{6}}=-\frac{17}{91\,440}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4x)}{\ln(\tan x)}\,dx+\frac{1}{1\,524}\int_{0}^{\frac{\pi}{4}}\frac{\sin(8x)}{\ln(\tan x)}\,dx-\frac{1}{2\,032}\int_{0}^{\frac{\pi}{4}}\frac{\sin(12x)}{\ln(\tan x)}\,dx\quad(1.3)$
These results are highlighting general patterns. In fact, it is easy to recognize the pattern
@@
Deeper linear algebra through Gauss elimination algorithm on (1.1), (1.2) and (1.3) allows to conclude following :
@@
After some resilience and tact, the following equalities also hold true :
@@
On this basis, we are led to state the following conjecture :
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(4nx)}{\ln(\tan x)}\,dx=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}$
The proof of this statement is fully detailed in [3] and is omitted here in the interest of brevity. We merely note the existence of integrals $I_{n}$ satisfying
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}\quad(2.4)$
Furthermore, we also find in [3] some integrals $I_{n}$ which adhere to this form
$\forall n\in\mathbb{N}^{*},\exists\,\left\{C_{p,n}:p\in\llbracket 1,n\rrbracket\right\}\subset\mathbb{Q}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\beta(2p)}{\pi^{2p-1}}\quad(2.5)$
The formal existence of integrals of the type (2.4) or (2.5) can already be traced back to the work of Tobias Kyrion [11], where the author investigates, among others, the following integrals:
@@
These three families of integrals were later rediscovered in [3] within a broader analytic framework, and are adopted here in a different formulation that is more suitable for the evaluation method and for discussing certain linear independence conjectures. Moreover, several additional integrals also satisfy either (2.4) or (2.5). The collection of all such integrals, together with the remarkable structural patterns they exhibit, constitutes the main contribution of this article. A representative selection is summarized in the following table, where $n$ denotes a strictly positive integer and $C_{p,n}$ are rational coefficients with $p\in\llbracket 1,n\rrbracket$.
@@
Proof
Let $J_{n}:=\displaystyle\int_{0}^{1}\dfrac{\mathrm{Li}_{-2n-1}(-x^{2})\ln\ln\frac{1}{x}}{x}\,dx$. On recalling the definition of the polylogarithm $\mathrm{Li}_{s}(z):=\sum_{k=1}^{\infty}\dfrac{z^{k}}{k^{s}}$ and bearing in mind that $|-x^{2}|<1$ on the domain of integration, one becomes
@@
After switching the order of the operators (assumed as justified), one has
@@
The later sum is then split in two to provide
@@
Let $\chi(s):=\left(\dfrac{\pi}{2}\right)^{-s}\sin\left(\dfrac{\pi s}{2}\right)\Gamma(s)$, so we get $\beta(1-s)=\chi(s)\,\beta(s)$. On differentiating with respect to $s$, we get
@@
. Let $\Delta(s):=\left(\dfrac{\pi}{2}\right)^{-s}\Gamma(s)$, so that $\chi(s)=\Delta(s)\sin\left(\dfrac{\pi s}{2}\right)$. On differentiating with respect to $s$, we have
@@
On evaluating this equality for $s=2n$ and putting it back in the earlier relation, we obtain
$\beta^{\prime}(1-2n)=(-1)^{n+1}\dfrac{2^{2n-1}(2n-1)!}{\pi^{2n-1}}\beta(2n)\quad(2.8)$
@@
The Malmsten’s integral becomes
@@
The interchange of summation and integration being assumed as correct. On evaluating the integral, we get
@@
One reads easily from the functional equation of $\beta$ that $\beta(1-2n)=0$. The value of $\beta^{\prime}(1-2n)$ follows from (2.8).
@@
with rational coefficients. Derivations steps detailed in [3] and omitted here for the sake of brevity. The following lines only give instructions on how to prove that
@@
holds form (2.4) and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ holds form (2.5)
The study of the integrals $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ relays solely on contour integration techniques by bearing in mind first of all the evenness of the integrands which imply that the integration bounds can be extended to $\infty$$\infty$[. We made the remark that the Laurent expansion approach is very fruitful to find the residues here. To find these residues, we note that $\cosh^{n}\!z$ has a pole of $n$-th order at $z_{l}:=\dfrac{(2l+1)i\pi}{2}$ where $l\in\mathbb{Z}$. We find the Laurent expansion of $\dfrac{1}{\cosh^{n}\!z}{}$ around $z_{l}$ by letting first of all $z=z_{l}+w$. It comes
@@
since $\cosh z_{l}=0$. Hence, finding the Taylor expansion of $\cosh^{n}\!z$ around $z_{l}$ reduces to finding the Taylor expansion of $\sinh^{n}z_{l}\,\sinh^{n}w$ around $w=0$, which is done after linearizing $\sinh^{n}w$ depending of the parity of $n$ and using the Taylor expansion of $\sinh(\alpha w)$ around $w=0$. With the Cauchy’s product formula we exhibit the Laurent expansion of $\cosh^{-n}\!z$ around $z_{l}$ and remark rational coefficients appearing in these series. On noticing additionally that $\dfrac{\sinh\!\left((2k+1)z\right)}{z}$ is analytic around $z=z_{l}$, we find its Taylor expansion by using the Leibniz’s formula for derivatives of products on bearing in mind these formulae which are quite easy to validate through induction on $n$:
@@
The proofs of Malmsten’s integrals outlined in Propositions 2.6 and 2.7 lack full mathematical rigor for two main reasons. First, the interchange of summation and integration is not properly justified. Second, the manipulations involve regularized versions of the beta and zeta functions, whose corresponding real-valued expressions would otherwise diverge. To establish a rigorous proof of these formulae, it is therefore necessary to derive explicit expressions for the coefficients $L_{p,n}$ and $K_{p,n}$ satisfying the given relations :
@@
This method is fully explored and detailled in [3]. It starts by showing first of all the following identities :
@@
In virtue of Proposition (2.6), the irrationality of $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ is equivalent to this conjecture :
@@
In virtue of Proposition (2.7), the irrationality of $\dfrac{\beta(2n)}{\pi^{2n}}$ is equivalent to this conjecture :
@@
Proof :
One has:
$I_{1}=x_{1,1}J_{1}$
$I_{2}=x_{1,2}J_{1}+x_{2,2}J_{2}$
$\vdots$
$I_{n}=x_{1,n}J_{1}+x_{2,n}J_{2}+\ldots+x_{n,n}J_{n}$
The last exercise consists of expressing each $J_{k}$ as linear combination of $I_{k}$ which corresponds to the last step of Gauss elimination algorithm.
Corollary : Since there exist integrals $I_{n}$ either holding form (2.4) or (2.5), we can always find rational coefficients $y_{k,n}$ satisfying $\widetilde{\beta\zeta}(n)=\displaystyle\sum_{k=1}^{n}y_{k,n}I_{k}$. The irrationality of $\dfrac{\widetilde{\beta\zeta}(n)}{\pi}$ would be deduced then from this conjecture :
@@
  • From the table of paragraph §2 and in combination with proposition 4.1, we conclude the existence of rational coefficients $\chi_{p,n}$ and $\xi_{p,n}$ which satisfy
@@
    The irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ would be deduced from this lonely following conjecture:
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $r_{p,n}$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}r_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Xi_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\int_{0}^{1}\frac{x\,\Xi_{n}(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx$.
@@
    On multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\beta(2n)}{\pi^{2n}}$ for each $n$ would follow from the following conjecture:
@@
  • On putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$, proposition 4.1 ensures the existence of some rational coefficients $s_{p,n}$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}s_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Lambda_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\int_{0}^{1}\frac{x\,\Lambda_{n}(x)}{\text{arctanh }x}\,dx$.
    Few first polynomials $\Lambda_{n}$ are [3]:
@@
    On multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ for each $n$ would follow from the following conjecture:
@@
Since proposition 4.1 allows to conclude that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\sum_{p=1}^{n}r_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4px)}{\ln(\tan x)}\,dx$ and that
@@
, with $r_{p,n},\Psi_{p,n}\in\mathbb{Q}$, the irrationality of the numbers $\dfrac{\beta(2n)}{\pi^{2n}}$ and $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ would be deduced from this conjecture :

### v3

@@
Abstract
We investigate special values of the Riemann zeta function at odd integers and of the Dirichlet beta function at even integers. We collect several analytic frameworks leading to these values and thereby provide a unifying perspective. Beyond their analytic interest, the resulting formulae motivate linear-independence conjectures which, if proved, would imply the irrationality of $\zeta(2n+1)/\pi^{2n+1}$ and $\beta(2n)/\pi^{2n}$.
@@ Introduction
Motivated by Euler’s evaluation of $\zeta(2n)$ and the expression of $\beta(2n+1)$, many authors tried to find some rational coefficients $r_{n}$ and $s_{n}$ such that the given relations may be satisfied
@@
Given the futility of the extraordinarily numerous attempts to compute these numbers, the solution of the designated problem is generally regarded as impossible; but so far a strong proof of this impossibility is lacking. Only heuristic methods and deep computations by brute force support this impossibility to this day. One may solve the problem by proving that the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ are not multiples of $\pi$. The difficulty of this new reformulation partly stems from the lack of sufficiently many closed-form expressions or identities involving these values that might serve as a starting point for an irrationality or transcendence proof. This paper is devoted to some integral representations of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n}}$ and $\dfrac{\beta(2n)}{\pi^{2n-1}}$ through interesting logarithmic, trigonometric and hyperbolic kernels, along with several conjectures that support their non-representability as rational multiples of $\pi$.
@@
that was first evaluated by Iaroslav V, Blagouchine in 2014 [3] and shown to equal $\dfrac{7\zeta(3)}{8\pi^{2}}$. Therefore the proof of its convergence is omitted here for the sake of conciseness. Notwithstanding, the presence of singularities at its bounds requires some care. Consequently, we rewrite the integral as the limit $I=\displaystyle\lim_{\xi\to 0^{+}}\int_{\xi}^{1-\xi}\dfrac{x(x^{4}-4x^{2}+1)\ln\ln\frac{1}{x}}{(1+x^{2})^{4}}dx$
@@
$I=\displaystyle\lim_{\xi\to 0^{+}}\dfrac{1}{2}\int_{\xi}^{1-\xi}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}dx=\dfrac{1}{2}\int_{0}^{1}\dfrac{x(x^{2}-1)}{(1+x^{2})^{3}\ln x}\,dx$
The change of variable $x\rightarrow\tan u$ yields:
@@
This gives finally
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4u}{\ln(\tan u)}\,du=-\dfrac{7\zeta(3)}{\pi^{2}}$ (1)
@@
$\dfrac{\zeta(5)}{\pi^{4}}=-\dfrac{1}{186}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 4x}{\ln(\tan x)}\,dx+\dfrac{1}{124}\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 8x}{\ln(\tan x)}\,dx$ (2)
$\frac{\zeta(7)}{\pi^{6}}=-\frac{17}{91\,440}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4x)}{\ln(\tan x)}\,dx+\frac{1}{1\,524}\int_{0}^{\frac{\pi}{4}}\frac{\sin(8x)}{\ln(\tan x)}\,dx-\frac{1}{2\,032}\int_{0}^{\frac{\pi}{4}}\frac{\sin(12x)}{\ln(\tan x)}\,dx$ (3)
These results highlight general patterns. In fact, it is easy to recognize the pattern
@@
Deeper linear algebra through the Gauss elimination algorithm on (1), (2) and (3) allows us to conclude the following :
@@
After further computations, the following equalities also hold true :
@@
[…]
$+8\,387\,584\dfrac{\zeta(13)}{\pi^{12}}$
$\int_{0}^{\frac{\pi}{4}}\dfrac{\sin 28x}{\ln(\tan x)}\,dx=\;$ $-\dfrac{88\,069}{45\,045}\dfrac{\zeta(3)}{\pi^{2}}+\dfrac{1\,409\,632}{7\,425}\dfrac{\zeta(5)}{\pi^{4}}-\dfrac{8\,091\,424}{945}\dfrac{\zeta(7)}{\pi^{6}}$
$+\dfrac{30\,685\,696}{135}\dfrac{\zeta(9)}{\pi^{8}}-3\,668\,224\dfrac{\zeta(11)}{\pi^{10}}+33\,550\,336\dfrac{\zeta(13)}{\pi^{12}}$
$-134\,213\,632\dfrac{\zeta(15)}{\pi^{14}}$
[…]
$-\dfrac{1\,878\,990\,848}{3}\dfrac{\zeta(15)}{\pi^{14}}+2\,147\,467\,264\dfrac{\zeta(17)}{\pi^{16}}$
These considerations motivate the following conjecture:
$\forall n\in\mathbb{N}^{*},\exists\,\begin{pmatrix}C_{1,n}\\ C_{2,n}\\ \vdots\\ C_{n,n}\end{pmatrix}\in\mathbb{Q}^{n}\quad\text{such that}\quad\int_{0}^{\frac{\pi}{4}}\dfrac{\sin(4nx)}{\ln(\tan x)}\,dx=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}$
The proof of this statement is fully detailed in [10, 59-67] and is omitted here in the interest of brevity. We merely note the existence of integrals $I_{n}$ satisfying
$\forall n\in\mathbb{N}^{*},\exists\,\begin{pmatrix}C_{1,n}\\ C_{2,n}\\ \vdots\\ C_{n,n}\end{pmatrix}\in\mathbb{Q}^{n}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\zeta(2p+1)}{\pi^{2p}}$ (4)
Furthermore, we also find in [10, 49-86] some integrals $I_{n}$ which adhere to this form
$\forall n\in\mathbb{N}^{*},\exists\,\begin{pmatrix}C_{1,n}\\ C_{2,n}\\ \vdots\\ C_{n,n}\end{pmatrix}\in\mathbb{Q}^{n}\quad\text{such that}\quad I_{n}=\sum_{p=1}^{n}C_{p,n}\,\frac{\beta(2p)}{\pi^{2p-1}}$ (5)
The formal existence of integrals of the type (4) or (5) can already be traced back to the work of Tobias Kyrion [6], where the author investigates, among others, the following integrals:
@@
These three families of integrals were later rediscovered in [10] within a broader analytic framework, and are adopted here in a different formulation that is more suitable for the evaluation method and for discussing certain linear independence conjectures. Moreover, several additional integrals also satisfy either (4) or (5). The collection of all such integrals, together with the remarkable structural patterns they exhibit, constitutes the main contribution of this article. A representative selection is summarized in the following table, where $n$ denotes a strictly positive integer and $C_{p,n}$ are rational coefficients with $p\in\llbracket 1,n\rrbracket$.
@@
Proof.
Let $J_{n}:=\displaystyle\int_{0}^{1}\dfrac{\mathrm{Li}_{-2n-1}(-x^{2})\ln\ln\frac{1}{x}}{x}\,dx$. Recalling the definition of the polylogarithm $\mathrm{Li}_{s}(z):=\sum_{k=1}^{\infty}\dfrac{z^{k}}{k^{s}}$ and bearing in mind that $|-x^{2}|<1$ on the domain of integration, one gets
@@
After switching the order of the operators (under suitable regularization assumptions), one has
@@
The latter sum is then split in two to provide
@@
Let $\chi(s):=\left(\dfrac{\pi}{2}\right)^{-s}\sin\left(\dfrac{\pi s}{2}\right)\Gamma(s)$, so we get $\beta(1-s)=\chi(s)\,\beta(s)$. Differentiating with respect to $s$, we get
@@
. Let $\Delta(s):=\left(\dfrac{\pi}{2}\right)^{-s}\Gamma(s)$, so that $\chi(s)=\Delta(s)\sin\left(\dfrac{\pi s}{2}\right)$. Differentiating with respect to $s$, we have
@@
Evaluating this equality for $s=2n$ and putting it back in the earlier relation, we obtain
$\beta^{\prime}(1-2n)=(-1)^{n+1}\dfrac{2^{2n-1}(2n-1)!}{\pi^{2n-1}}\beta(2n)$ (6)
@@
The Malmsten’s integral becomes this after interchanging summation and integration (under suitable convergence assumptions)
@@
After evaluating the integral, we get
@@
One reads easily from the functional equation of $\beta$ that $\beta(1-2n)=0$. The value of $\beta^{\prime}(1-2n)$ follows from (6).
@@
with rational coefficients. The derivations are detailed in [10] and omitted here for the sake of brevity. The following lines only give instructions on how to prove that
@@
is of the form (4) and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ is of the form (5)
The study of the integrals $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n}\!x}\,dx$ and $\int_{0}^{+\infty}\frac{\sinh\!\left((2k+1)x\right)}{x\cosh^{2n+1}\!x}\,dx$ relies solely on contour integration techniques by bearing in mind first of all the evenness of the integrands which imply that the integration bounds can be extended to $\left(-\infty,+\infty\right)$. We made the remark that the Laurent expansion approach is very fruitful to find the residues here. To find these residues, we note that $\cosh^{n}\!z$ has a pole of $n$-th order at $z_{l}:=\dfrac{(2l+1)i\pi}{2}$ where $l\in\mathbb{Z}$. We find the Laurent expansion of $\dfrac{1}{\cosh^{n}\!z}{}$ around $z_{l}$ by letting first of all $z=z_{l}+w$. It comes
@@
since $\cosh z_{l}=0$. Hence, finding the Taylor expansion of $\cosh^{n}\!z$ around $z_{l}$ reduces to finding the Taylor expansion of $\sinh^{n}z_{l}\,\sinh^{n}w$ around $w=0$, which is done after linearizing $\sinh^{n}w$ depending of the parity of $n$ and using the Taylor expansion of $\sinh(\alpha w)$ around $w=0$. With the Cauchy’s product formula we exhibit the Laurent expansion of $\cosh^{-n}\!z$ around $z_{l}$ and remark rational coefficients appearing in these series. Since $\dfrac{\sinh\!\big((2k+1)z\big)}{z}$ is analytic near $z=z_{l}$, its Taylor expansion follows from Leibniz’s rule for derivatives of products together with the following identities, which are readily proved by induction on $n$.
@@
The proofs of Malmsten’s integrals outlined in proposition˜2.1 and proposition˜2.2 lack full mathematical rigor for two main reasons. First, the interchange of summation and integration is not properly justified. Second, the manipulations involve regularized versions of the beta and zeta functions, whose corresponding real-valued expressions would otherwise diverge. To establish a rigorous proof of these formulae, it is therefore necessary to derive explicit expressions for the coefficients $L_{p,n}$ and $K_{p,n}$ satisfying the given relations :
@@
This method is fully explored and detailed in [10, 86-93, 123-130]. It starts by showing first of all the following identities :
@@
By virtue of proposition˜2.1, the irrationality of $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ is equivalent to this conjecture :
@@
By virtue of proposition˜2.2, the irrationality of $\dfrac{\beta(2n)}{\pi^{2n}}$ is equivalent to this conjecture :
@@
Proof.
The exercise consists of expressing each $J_{k}$ as linear combination of $I_{k}$ which corresponds to the last step of Gauss elimination algorithm. The result is another formulation of the following statement :"The inverse of a lower triangular matrix is again a lower triangular matrix."∎
Corollary : Since there exist integrals $I_{n}$ either holding form (4) or (5), we can always find rational coefficients $y_{k,n}$ satisfying $\widetilde{\beta\zeta}(n)=\displaystyle\sum_{k=1}^{n}y_{k,n}I_{k}$. The irrationality of $\dfrac{\widetilde{\beta\zeta}(n)}{\pi}$ would follow then from this conjecture :
@@
  • From the table of section˜2 and in combination with proposition˜4.1, we conclude the existence of rational coefficients $\chi_{p,n}$ and $\xi_{p,n}$ which satisfy
@@
    The irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ and $\dfrac{\beta(2n)}{\pi^{2n}}$ would follow from this single conjecture:
@@
  • Putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$, proposition˜4.1 ensures the existence of some rational coefficients $r_{p,n}$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\sum_{p=1}^{n}r_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\sqrt{1-x^{2}}\,\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Xi_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\beta(2n)}{\pi^{2n-1}}=\displaystyle\int_{0}^{1}\frac{x\,\Xi_{n}(x)}{\sqrt{1-x^{2}}\,\text{arctanh }x}\,dx$.
@@
    After multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\beta(2n)}{\pi^{2n}}$ for each $n$ would follow from the following conjecture:
@@
  • Putting the focus on the integral $\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$, proposition˜4.1 ensures the existence of some rational coefficients $s_{p,n}$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\sum_{p=1}^{n}s_{p,n}\displaystyle\int_{0}^{1}\frac{x^{2p-1}}{\operatorname{arctanh}\,x}\,dx$; thus, the existence of polynomials $\Lambda_{n}\in\mathbb{Q}[x]$ such that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\displaystyle\int_{0}^{1}\frac{x\,\Lambda_{n}(x)}{\text{arctanh }x}\,dx$.
    Few first polynomials $\Lambda_{n}$ are [10, 131-132]:
@@
    After multiplying each polynomial in $\mathbb{Q}[x]$ by the lowest common multiple of denominators of its coefficients, we get a polynomial in $\mathbb{Z}[x]$. Hence, the irrationality of the numbers $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$ for each $n$ would follow from the following conjecture:
@@
Since proposition˜4.1 shows that $\dfrac{\zeta(2n+1)}{\pi^{2n}}=\sum_{p=1}^{n}r_{p,n}\int_{0}^{\frac{\pi}{4}}\frac{\sin(4px)}{\ln(\tan x)}\,dx$ and that
@@
, with $r_{p,n},\Psi_{p,n}\in\mathbb{Q}$, this conjecture, if true, would provide a pathway toward establishing the irrationality of $\dfrac{\beta(2n)}{\pi^{2n}}$ and $\dfrac{\zeta(2n+1)}{\pi^{2n+1}}$: